OpenStudy (anonymous):

A rectangular pyramid has a volume of 756 cubic feet. If the base has a length of 9 feet and a width of 14 feet, what is the height of the pyramid?

OpenStudy (saifoo.khan):

Do you know the formula for the volume of a pyramid?

OpenStudy (anonymous):

yeah v =1/3bh

OpenStudy (saifoo.khan):

Now, insert the given values one by one in it.

OpenStudy (anonymous):

1/3 lwh = v

OpenStudy (saifoo.khan):

it's http://chemwiki.ucdavis.edu/ @api/deki/files/4758/=Volume_of_a_Pyramid.jpg

OpenStudy (anonymous):

plz no links i cant open the links

Directrix (directrix):

V = (1/3) B h where B is area of the Base and h is the height of the pyramid. 756 = (1/3) 9 (14) h --> 9*14 is area of rectangular base 756 = 3 *14 * h 756/(3*14) = h 756/42 = h 18 = h ----> Check my work.

OpenStudy (anonymous):

thx at first iwasnt sure if it was 18 or 6

OpenStudy (anonymous):

i think i used the wrong formula, thx alot